OpenStudy (anonymous):

Solve for the variable in the following proportion. 1/4 is to 1 1/4 as 2 is to b b=__?

OpenStudy (anonymous):

(1/4) / (1 1/4) = 2 / b Calculating LHS: 0.2 = 2/b b = 2/0.2 b = ?

OpenStudy (anonymous):

10?